Given the following measurements from John’s health assessment, answer the question given:

            End diastolic volume = 105 ml

            End systolic volume = 30 ml

            Heart rate = 70 bpm

John’s maximum heart rate was calculated to be 200 bpm. What is the appropriate training range for John?

(0.1pts)

Question 83 - Given the following measurements from John’s health assessment, answer the question given:             End diastolic volume = 105 ml             End systolic volume = 30 ml             Heart rate = 70 bpm John’s maximum heart rate was calculated to be 200 bpm. What is the appropriate training range for John?

Select 148 – 174 bpm as your answer

148 – 174 bpm

Select 120 – 180 bpm as your answer

120 – 180 bpm

Select 78 – 104 bpm as your answer

78 – 104 bpm

Select 190 – 250 bpm as your answer

190 – 250 bpm
Answer #1

the answer is 120-180 bpm

SOLUTION -

  • A heart rate training zone is a range that defines the upper and lower limits of training intensities. It is calculated using an age-related predicted maximum heart rate (HRmax) and a special equation called heart rate reserve .
  • John's predicted age is around 20 years as per his maximum heart rate.
  • The Heart Rate Reseve Equation Determines the Training Range as per the following steps:-

1. Maximum Heart Rate (given) - Age = HRmax

  2. HRmax - Resting Heart Rate = Heart Rate Reserve (HRR)

3. HRR X percentage that JOHN wants to train at = Z

4. Z + Resting Heart Rate = Training Heart Rate

Note- Percentage at JOHN wants to train at = (Maximum Heart Rate - Resting Heart Rate ) / Maximum Heart Rate/ 100

= 200- 70/200/100

= 0.65

so as per the above steps

1. 200 - 20 = 180

2. 180 - 70 = 110

3. 110 X 0.65 = 71.50

4. 71.50 + 70 = 141.50

SO THE APPROPRIATE TRAINING RANGE FOR JOHN IS 120-180 bpm